# Making Images Real Again: A Comprehensive Survey on Deep Image Composition

### v7

@@ Abstract
As a common image editing operation, image composition (object insertion) aims to combine the foreground from one image and another background image, resulting in a composite image. However, there are many issues that could make the composite images unrealistic. These issues can be summarized as the inconsistency between foreground and background, which includes appearance inconsistency (e.g., incompatible illumination), geometry inconsistency (e.g., unreasonable size), and semantic inconsistency (e.g., mismatched semantic context). Image composition task could be decomposed into multiple sub-tasks, in which each sub-task targets at one or more issues. Specifically, object placement aims to find reasonable scale, location, and shape for the foreground. Image blending aims to address the unnatural boundary between foreground and background. Image harmonization aims to adjust the illumination statistics of foreground. Shadow (resp., reflection) generation aims to generate plausible shadow (resp., reflection) for the foreground. These sub-tasks can be executed sequentially or parallelly to acquire realistic composite images. To the best of our knowledge, there is no previous survey on image composition (object insertion). In this paper, we conduct comprehensive survey over the sub-tasks and combinatorial task of image composition (object insertion). For each one, we summarize the existing methods, available datasets, and common evaluation metrics. Datasets and codes for image composition are summarized at https://github.com/bcmi/Awesome-Image-Composition. We have also contributed the first image composition toolbox: libcom https://github.com/bcmi/libcom, which assembles 10+ image composition related functions (e.g., image blending, image harmonization, object placement, shadow generation, generative composition). The ultimate goal of this toolbox is solving all the problems related to image composition with simple ‘import libcom’.
@@ I Introduction
Figure 2: The quality of composite image is degraded by the appearance inconsistency, geometric inconsistency, and semantic inconsistency. Each type of inconsistency involves a number of issues. Each sub-task targets at addressing one or more issues.
Figure 3: Previous works perform multiple sub-tasks (e.g., object placement, image blending, image harmonization, shadow/reflection generation) sequentially or parallelly to achieve the goal of image composition.
TABLE I: The issues to be solved in image composition task and the corresponding deep learning methods to solve these issue. Note that some methods only focus on one issue while some methods attempt to solve multiple issues simultaneously. “Boundary” means refining the boundary between foreground and background. “Appearance” means adjusting the illumination of foreground. “Shadow” means generating shadow for the foreground. “Reflection” means generating reflection for the foreground. “Geometry” means seeking for reasonable location, size, and shape for the foreground considering geometric constraints. “Occlusion” means coping with the unreasonable occlusion. “Semantics” means finding suitable semantic context for the object.
@@
The appearance inconsistency is including but not limited to: 1) unnatural boundary between foreground and background; 2) incompatible illumination statistics between foreground and background; 3) missing or implausible shadow and reflection of foreground; 4) resolution, sharpness, and noise discrepancy between foreground and background [94]. For the first issue, the foreground is usually extracted using image segmentation [129] or matting [218, 42] algorithms. However, the foregrounds may not be precisely delineated, especially at the boundaries. When pasting the foreground with jagged boundaries on the background, there would be obvious color artifacts along the boundary. To solve this issue, image blending [210, 241] aims to address the unnatural boundary between foreground and background, so that the foreground could be seamlessly blended with the background. For the second issue, since the foreground and background may be captured in different conditions (e.g., weather, season, time of the day, camera setting), the obtained composite image could be inharmonious (e.g., foreground captured in the daytime and background captured at night). To solve this issue, image harmonization [185, 29, 25] aims to adjust the illumination statistics of foreground to make it more compatible with the background, so that the whole image looks more harmonious. For the third issue, when pasting the foreground on the background, the foreground may also affect the background with shadow or reflection. To solve this issue, shadow generation [111, 248, 158] or reflection generation [125, 208, 180] focus on generating plausible shadow or reflection for the foreground according to both foreground and background information. For the fourth issue, the foreground and background may be from two images with different resolutions, blur degrees, and noise patterns. The resolution (resp., sharpness, noise) discrepancy between them could be mitigated by using super-resolution [204], deblurring [239], denoising [182] techniques.
@@
The geometric inconsistency is including but not limited to: 1) the foreground object is too large or too small; 2) the foreground object does not have reasonable supporting force (e.g., hanging in the air); 3) unreasonable occlusion; 4) inconsistent perspectives between foreground and background. In summary, the location, size, and shape of the foreground may be irrational considering the geometric constraints. Object placement [3, 33, 78, 183, 240] tends to seek for reasonable location, size, and shape by predicting the foreground transformation to avoid the abovementioned inconsistencies. Previous object placement methods [240, 183] mainly predict simple form of spatial transformation, that is, shifting and scaling the foreground to achieve reasonable location and size. Some other methods [78, 107] predict more general form of spatial transformation (e.g., affine transformation, perspective transformation, thin plate spline transformation) to warp the foreground. In terms of more advanced geometric transformation like view synthesis and pose transfer, we should resort to generative approaches [222, 167] to change the viewpoint/pose of the foreground. When placing the object on the background, unreasonable occlusion may occur. Most previous methods seek for reasonable placement to avoid unreasonable occlusions, while some methods [3, 233, 174] aim to fix unreasonable occlusion by removing the occluded regions of foreground based on the estimated depth information.
@@
The semantic inconsistency is including but not limited to: 1) the foreground appears at a semantically unreasonable place (e.g., a zebra is placed in the living room); 2) the foreground has unreasonable interactions with other objects or people (e.g., a person is riding a motorbike, but the person and the motorbike are facing towards opposite directions); 3) the background may have semantic impact on the foreground appearance. The semantic inconsistency is judged based on commonsense knowledge, so the cases of semantic inconsistency may be arguable according to subjective judgement. For example, when a car is placed in the water, it can be argued that a car is sinking into the water after a car accident. However, such event has rather low probability compared with commonly seen cases, so we can claim that the car appears at an unreasonable place, which belongs to semantic inconsistency. Partial solution to semantic inconsistency falls into the scope of object placement. To be exact, by predicting suitable spatial transformation for the foreground, we can relocate the foreground to a reasonable place or adjust the pose of foreground to make its interactions with environment more convincing. Additionally, the appearance of foreground object may be affected by the background semantically, which is different from low-level appearance inconsistency (illumination, shadow). For example, a car placed on the snowy ground may be covered by snow. A student inserted into a group of students wearing school uniforms should wear the same school uniform. Such semantic appearance variation is very flexible and challenging, which will not be fully discussed in this survey.
@@
So far, we have introduced several sub-tasks (e.g., image blending, image harmonization, shadow generation, object placement), in which one sub-task targets at one or multiple issues. Previous works usually focus on one sub-task or perform multiple sub-tasks sequentially (i.e., image blending followed by image harmonization) as shown in Fig. 3(a). The reasonable sequential order is as follows. Given a pair of foreground and background, we first use object placement to find suitable scale and location for the foreground, and use image blending to refine the boundary between foreground and background. Then, we use image harmonization to adjust the foreground illumination and shadow generation to generate plausible shadow for the foreground. Recently, as the diffusion models have demonstrated unprecedented generation ability, some works [222, 167] utilize diffusion models to perform multiple sub-tasks (e.g., image blending, image harmonization, view synthesis) parallelly as shown in Fig. 3(b). Given a pair of foreground and background with bounding box, they propose one unified model to directly produce the composite image, in which the foreground is blended seamlessly and harmoniously into the background. These methods re-generate the foreground object instead of making restrained adjustments for the foreground object, so we refer to them as generative image composition methods. We summarize all the potential issues and the corresponding methods to solve them in Table I. Note that we use $\cdot$ instead of + in some slots, because these issues are only partially solved. Another problem for parallel pipeline is that whether we can switch on or switch off some subtasks when performing them parallelly, with the goal of more controllable image composition. Based on this idea, Zhang et al. [234] built a prototype named ControlCom, which uses a 2-dim binary indicator vector to selectively adjust the illumination or pose of foreground object. In the ideal case, assuming that we have in total $K$ subtasks, a $K$-dim binary indicator vector could be used to indicate the selected subtasks.
@@
In the remainder of this paper, we will elaborate on each sub-task or combinatorial task. In particular, we will introduce object placement in Section II, image blending in Section III, image harmonization in Section IV, shadow (resp., reflection) generation in Section V (resp., VI), generative image composition in Section VII, foreground object search in Section VIII. In each section, we will introduce the existing methods, available datasets, and common evaluation metrics. Finally, we will conclude the whole paper in Section IX. The contributions of this paper can be summarized as follows,
@@
  • For each sub-task and combinatorial task, we survey the existing methods, available datasets, and common evaluation metrics. We believe that this comprehensive survey can serve as the roadmap for the future research in a broad community.
@@ II Object Placement
Object placement aims to paste the foreground on the background with suitable location, size, and shape. As shown in Fig. 4, the cases of unreasonable object placement are including but not limited to: a) the foreground object has inappropriate size (e.g., the dog is too large); b) the foreground object has unreasonable occlusion with background objects (e.g., the fences are unreasonably occluded by the giraffe); c) the foreground object does not have reasonable force condition (e.g., the suitcase is floating in the air); d) the foreground object appears at a semantically unreasonable place (e.g., the boat appears on the land); e) inconsistent perspectives between foreground and background (e.g., the car and the bus have inconsistent perspectives). By taking all the above factors into consideration, object placement is a very challenging task.
@@ II-B2 Instance-specific Object Placement
Generative approaches: Generative approaches [183, 240, 107, 100, 99, 224, 200] predict different types of spatial transformations (e.g., shifting and scaling, affine transformation, perspective transformation) for the foreground object, which is more flexible and powerful than category-specific object placement methods. For instance, Tripathi et al. [183] developed a model with generator, discriminator, and target network. Given a pair of background and foreground, the generator predicts the affine transformation for the foreground object to produce a composite image. The produced composite image is expected to fool the discriminator and fit the target network corresponding to a downstream task (e.g., object detection). Similarly, Zhan et al. [231] adopted spatial transformer network (STN) [70] to predict the warping parameters under an adversarial learning framework. To produce multiple reasonable placements, Zhang et al. [240] combined the foreground feature, background feature, and a random vector to predict the object placement. Moreover, they ensured the diversity of object placement by enforcing the pairwise distances between predicted placements to approach those between corresponding random vectors. To promote the diversity of generated placements, Zhou et al. [259] established the bijection between random vector and positive composite image. Moreover, they reformulated object placement as a graph completion task. In particular, background nodes have both content features and placements, while the inserted foreground node only has content feature, giving rise to an incomplete graph. Hence, they estimated the missing placement for the foreground node to complete the graph. Zhang et al. [246] proposed to make sequential decisions to produce a reasonable placement by using reinforcement learning. Qin et al. [146] employed a pre-trained large multi-modal model to generate a caption containing the placement information, and then predicted the placement bounding boxes. Zhou et al. [255] first predicted candidate bounding boxes and then associated each object query with matched bounding box.
@@
Discriminative approaches: Liu et al. [113] proposed a discriminative approach named SimOPA to verify whether a composite image is rational in terms of the foreground object placement. Particularly, they feed the concatenation of composite image and foreground mask into a binary classification network to predict a rationality score. However, this discriminative approach is very inefficient, because they need to go through the discriminative network multiple times to find a reasonable object placement. To address this issue, Niu et al. [132] dubbed SimOPA as slow object placement assessment (SOPA) model and proposed a fast object placement assessment (FOPA) model, which can predict the rationality scores at all locations by going through the model only once. Precisely, they take in a pair of background and scaled foreground, and produce a rationality score map, in which each entry represents the rationality score of the composite image obtained by pasting the foreground at this location. They developed several innovations (e.g., background prior transfer, feature mimicking) to bridge the performance gap between FOPA and SOPA, reaching the conclusion that FOPA can achieve comparable performance with SimOPA at significantly reduced cost. FOPA [132] has also demonstrated stronger ability to generate realistic composite images than generative approaches [259, 246]. Similar to FOPA [132], Zhu et al. [262] proposed to predict the rationality scores of all scales and locations, based on the interaction output between foreground and background using transformer [187]. Zhu et al. [262] also explored using unlabeled images with deliberately designed loss functions for object placement task.
In the end, we briefly discuss the occlusion issue. Most of the above methods seek for reasonable placements to avoid the occurrence of occlusion, i.e., the inserted foreground is not occluded by background objects. Differently, a few methods [3, 233, 174] attempt to address the unreasonable occlusion when it occurs. Specifically, they first estimate the relative depth relation between the foreground object and the surrounding background objects. Then, they remove the occluded part of foreground object. In this way, they are able to generate composite images with reasonable inter-object occlusions.
@@ III-D Experiments
By taking LFPNet [114] as an example matting method, we predict the alpha mattes and obtain the composite images. We observe that LFPNet can generally achieve satisfactory results except some challenging cases. We pick out its several failure cases to verify the effectiveness of image blending methods.
We report the results of Poisson image blending [143], GP-GAN [210], Zhang et al. [241], and MLF [237]. We also report the ground-truth composite image obtained using ground-truth alpha matte for comparison. From Fig. 9, it can be seen that the obtained composite images using predicted alpha mattes are very close to the ground-truth composite image except partial boundary regions. We observe that Poisson image blending [143] smooths the transition boundary to some extent, but unexpectedly distorts the foreground content by seeping through the foreground. GP-GAN [210] and Zhang et al. [241] are inspired by Poisson image blending, but use content loss to preserve the original foreground content. Therefore, they strike a balance between preserving the foreground content and smoothing the boundary. However, some smoothed boundary regions are still not satisfactory. MLF [237] can obtain visually appealing results in some cases. Nonetheless, it may erase detailed information (e.g., the small leaves of pineapple) and fail in handling transparent foreground objects (e.g., plastic bag).
For practical usage, traditional image blending methods are recommended under specific circumstances. For example, when the foreground mask is accurate and there is no high demand for sharp boundary, alpha blending [145] could be used. When the foreground has the same target color as surrounding background, Poisson blending [143] could be used. In other complex cases, learnable image blending methods [237, 234] are recommended, but they may struggle in dealing with transparent foreground and delicate edges. In the cases of transparent foreground (e.g., glass) and delicate edges (e.g., floating hair strands), it is suggested to improve the performance of image matting methods.

### v8

@@ Abstract
As a common image editing operation, image composition (object insertion) aims to combine the foreground from one image and another background image, to produce a composite image. However, there are many issues that could make the composite images unrealistic. These issues can be summarized as the inconsistency between foreground and background, which includes appearance inconsistency (e.g., incompatible illumination), geometry inconsistency (e.g., unreasonable size), and semantic inconsistency (e.g., mismatched semantic context). The image composition task could be decomposed into multiple sub-tasks, in which each sub-task targets one or more issues. Specifically, object placement aims to find reasonable scale, location, and shape for the foreground. Image blending aims to address the unnatural boundary between foreground and background. Image harmonization aims to adjust the illumination statistics of foreground. Shadow (resp., reflection) generation aims to generate plausible shadow (resp., reflection) for the foreground. These sub-tasks can be executed sequentially or in parallel to acquire realistic composite images. To the best of our knowledge, there is no previous survey on image composition. In this paper, we conduct a comprehensive survey over the sub-tasks and combined task of image composition. For each one, we summarize the existing methods, available datasets, and common evaluation metrics. Datasets and codes for image composition are summarized at https://github.com/bcmi/Awesome-Image-Composition. We have also contributed the first image composition toolbox: libcom https://github.com/bcmi/libcom, which assembles 10+ image-composition-related functions (e.g., image blending, image harmonization, object placement, shadow generation, generative composition). The ultimate goal of this toolbox is to solve all image composition problems with simple ‘import libcom’.
@@ I Introduction
Figure 2: The quality of composite image is degraded by the appearance inconsistency, geometric inconsistency, and semantic inconsistency. Each type of inconsistency involves a number of issues. Each sub-task targets one or more issues.
Figure 3: Previous works perform multiple sub-tasks (e.g., object placement, image blending, image harmonization, shadow/reflection generation) sequentially or in parallel to achieve the goal of image composition.
TABLE I: The issues to be solved in image composition task and the corresponding deep learning methods to solve these issue. Note that some methods only focus on one issue while some methods attempt to solve multiple issues simultaneously. “Boundary” means refining the boundary between foreground and background. “Appearance” means adjusting the illumination of foreground. “Shadow” means generating shadow for the foreground. “Reflection” means generating reflection for the foreground. “Geometry” means seeking reasonable location, size, and shape for the foreground considering geometric constraints. “Occlusion” means coping with the unreasonable occlusion. “Semantics” means finding suitable semantic context for the foreground.
@@
The appearance inconsistency includes, but is not limited to: 1) unnatural boundary between foreground and background; 2) incompatible illumination statistics between foreground and background; 3) missing or implausible shadow and reflection of foreground; 4) resolution, sharpness, and noise discrepancy between foreground and background [100]. For the first issue, the foreground is usually extracted using image segmentation [136] or matting [226, 45] algorithms. However, the foregrounds may not be precisely delineated, especially at the boundaries. When pasting the foreground with jagged boundaries on the background, there would be obvious color artifacts along the boundary. To solve this issue, image blending [218, 250] aims to address the unnatural boundary between foreground and background, so that the foreground could be seamlessly blended with the background. For the second issue, since the foreground and background may be captured in different conditions (e.g., weather, season, time of the day, camera setting), the obtained composite image could lack visual harmony (e.g., foreground captured in the daytime and background captured at night). To solve this issue, image harmonization [193, 32, 28] aims to adjust the illumination statistics of foreground to make it more compatible with the background, so that the resulting composite image looks more harmonious. For the third issue, when pasting the foreground on the background, the foreground may also affect the background with shadow or reflection. To solve this issue, shadow generation [118, 258, 166] or reflection generation [132, 216, 188] focus on generating plausible shadow or reflection for the foreground according to both foreground and background information. For the fourth issue, the foreground and background may be from two images with different resolutions, blur degrees, and noise patterns. The resolution (resp., sharpness, noise) discrepancy between them could be mitigated by using super-resolution [212], deblurring [248], denoising [190] techniques.
@@
The geometric inconsistency includes, but is not limited to: 1) the foreground object is too large or too small; 2) the foreground object does not have reasonable supporting force (e.g., hanging in the air); 3) unreasonable occlusion; 4) inconsistent perspectives between foreground and background. In summary, the location, size, and shape of the foreground may be irrational considering the geometric constraints. Object placement [3, 36, 83, 191, 249] tends to seek reasonable location, size, and shape by predicting the foreground transformation to avoid the abovementioned inconsistencies. Previous object placement methods [249, 191] mainly predict simple form of spatial transformation, that is, shifting and scaling the foreground to achieve reasonable location and size. Some other methods [83, 114] predict more general form of spatial transformation (e.g., affine transformation, perspective transformation, thin plate spline transformation) to warp the foreground. In terms of more advanced geometric transformation like view synthesis and pose transfer, we should resort to generative approaches [230, 175] to change the viewpoint/pose of the foreground. When placing the object on the background, unreasonable occlusion may occur. Most previous methods seek reasonable placement to avoid unreasonable occlusions, while some methods [3, 242, 182] aim to fix unreasonable occlusion by removing the occluded regions of foreground based on the estimated depth information.
@@
The semantic inconsistency includes, but is not limited to: 1) the foreground appears at a semantically unreasonable place (e.g., a zebra is placed in the living room); 2) the foreground has unreasonable interactions with other objects or people (e.g., a person is riding a motorbike, but the person and the motorbike are facing towards opposite directions); 3) the background may have semantic impact on the foreground appearance. The semantic inconsistency is judged based on commonsense knowledge, so the cases of semantic inconsistency may be arguable according to subjective judgement. For example, when a car is placed in the water, it can be argued that a car is sinking into the water after a car accident. However, such event has rather low probability compared with commonly seen cases, so we can claim that the car appears at an unreasonable place, which belongs to semantic inconsistency. Partial solution to semantic inconsistency falls into the scope of object placement. To be exact, by predicting suitable spatial transformation for the foreground, we can relocate the foreground to a reasonable place or adjust the pose of foreground to make its interactions with environment more convincing. Additionally, the appearance of foreground object may be affected by the background semantically, which is different from low-level appearance inconsistency (illumination, shadow). For example, a car placed on the snowy ground may be covered by snow. Another example is that a student inserted into a group of students wearing school uniforms should wear the same school uniform. Such semantic appearance variation is very flexible and challenging, which will not be fully discussed in this survey.
@@
So far, we have introduced several sub-tasks (e.g., image blending, image harmonization, shadow generation, object placement), in which one sub-task targets one or multiple issues. Previous works usually focus on one sub-task or perform multiple sub-tasks sequentially (i.e., image blending followed by image harmonization) as shown in Fig. 3(a). The reasonable sequential order is as follows. Given a pair of foreground and background, we first use object placement to find suitable scale and location for the foreground, and use image blending to refine the boundary between foreground and background. Then, we use image harmonization to adjust the foreground illumination and shadow generation to generate plausible shadow for the foreground. Recently, as the diffusion models have demonstrated unprecedented generation ability, some works [230, 175] utilize diffusion models to perform multiple sub-tasks (e.g., image blending, image harmonization, view synthesis) parallelly as shown in Fig. 3(b). Given a pair of foreground and background with bounding box, they propose one unified model to directly produce the composite image, in which the foreground is blended seamlessly and harmoniously into the background. These methods re-generate the foreground object instead of making restrained adjustments for the foreground object, so we refer to them as generative image composition methods. We summarize all the potential issues and the corresponding methods to solve them in Table I.
@@
In the remainder of this paper, we will elaborate on each sub-task or the combined task. In particular, we will introduce object placement in Section II, image blending in Section III, image harmonization in Section IV, shadow (resp., reflection) generation in Section V (resp., VI), generative image composition in Section VII, foreground object search in Section VIII. In each section, we will introduce the existing methods, available datasets, and common evaluation metrics. Finally, we will conclude the whole paper in Section IX. The contributions of this paper can be summarized as follows,
@@
  • For each sub-task and the combined task, we survey the existing methods, available datasets, and common evaluation metrics. We also show some experimental results. We believe that this comprehensive survey can serve as the roadmap for the future research in a broad community.
@@ II Object Placement
Object placement aims to paste the foreground on the background with suitable location, size, and shape. As shown in Fig. 4, the cases of unreasonable object placement include but are not limited to: a) the foreground object has inappropriate size (e.g., the dog is too large); b) the foreground object has unreasonable occlusion with background objects (e.g., the fences are unreasonably occluded by the giraffe); c) the foreground object does not have reasonable force condition (e.g., the suitcase is floating in the air); d) the foreground object appears at a semantically unreasonable place (e.g., the boat appears on the land); e) inconsistent perspectives between foreground and background (e.g., the car and the bus have inconsistent perspectives). By taking all the above factors into consideration, object placement is a very challenging task.
@@ II-B2 Instance-specific Object Placement
Generative approaches: Generative approaches [191, 249, 114, 106, 105, 233, 208, 27] predict different types of spatial transformations (e.g., shifting and scaling, affine transformation, perspective transformation) for the foreground object, which is more flexible and powerful than category-specific object placement methods. For instance, Tripathi et al. [191] developed a model with generator, discriminator, and target network. Given a pair of background and foreground, the generator predicts an affine transformation for the foreground object to produce a composite image. The produced composite image is expected to fool the discriminator and fit the target network corresponding to a downstream task (e.g., object detection). Similarly, Zhan et al. [240] adopted spatial transformer network (STN) [75] to predict the warping parameters under an adversarial learning framework. To produce multiple reasonable placements, Zhang et al. [249] combined the foreground feature, background feature, and a random vector to predict the object placement. Moreover, they ensured the diversity of object placement by enforcing the pairwise distances between predicted placements to approach those between corresponding random vectors. To promote the diversity of generated placements, Zhou et al. [270] established the bijection between random vector and positive composite image. Moreover, they reformulated object placement as a graph completion task. In particular, background nodes have both content features and placements, while the inserted foreground node only has content feature, giving rise to an incomplete graph. Hence, they estimated the missing placement for the foreground node to complete the graph. Zhang et al. [256] proposed to make sequential decisions to produce a reasonable placement by using reinforcement learning. Qin et al. [154] employed a pre-trained large multi-modal model to generate a caption containing the placement information, and then predicted the placement bounding boxes. Zhou et al. [266] first predicted candidate bounding boxes and then associated each object query with matched bounding box.
@@
Discriminative approaches: Liu et al. [120] proposed a discriminative approach named SimOPA to verify whether a composite image is rational in terms of the foreground object placement. Particularly, they feed the concatenation of composite image and foreground mask into a binary classification network to predict a rationality score. However, this discriminative approach is very inefficient, because they need to go through the discriminative network multiple times to find a reasonable object placement. To address this issue, Niu et al. [139] dubbed SimOPA as slow object placement assessment (SOPA) model and proposed a fast object placement assessment (FOPA) model, which can predict the rationality scores at all locations by going through the model only once. Precisely, they take in a pair of background and scaled foreground, and produce a rationality score map, in which each entry represents the rationality score of the composite image obtained by pasting the foreground at this location. They developed several innovations (e.g., background prior transfer, feature mimicking) to bridge the performance gap between FOPA and SOPA, reaching the conclusion that FOPA can achieve comparable performance with SimOPA at significantly reduced cost. FOPA [139] has also demonstrated stronger ability to generate realistic composite images than generative approaches [270, 256]. Similar to FOPA [139], Zhu et al. [273], Poska et al. [153] also proposed to predict the rationality scores of all scales and locations, based on the interaction output between foreground and background. Gao et al. [46] explored extending object placement task to a broader range of foreground categories and background scenes.
In the end, we briefly discuss the occlusion issue. Most of the above methods seek for reasonable placements to avoid the occurrence of occlusion, i.e., the inserted foreground is not occluded by background objects. Differently, a few methods [3, 242, 182, 54] attempt to address the unreasonable occlusion when it occurs. Specifically, they first estimate the relative depth relation between the foreground object and the surrounding background objects. Then, they remove the occluded part of foreground object. In this way, they are able to generate composite images with reasonable inter-object occlusions.
@@ III-D Experiments
By taking LFPNet [121] as an example matting method, we predict the alpha mattes and obtain the composite images. We observe that LFPNet can generally achieve satisfactory results except some challenging cases. We select several of its failure cases to verify the effectiveness of image blending methods.
We report the results of Poisson image blending [150], GP-GAN [218], Zhang et al. [250], and MLF [246]. We also report the ground-truth composite image obtained using ground-truth alpha matte for comparison. From Fig. 9, it can be seen that the composite images obtained using predicted alpha mattes are very close to the ground-truth composite image except partial boundary regions. We observe that Poisson image blending [150] smooths the transition boundary to some extent, but unexpectedly distorts the foreground content by seeping through the foreground. GP-GAN [218] and Zhang et al. [250] are inspired by Poisson image blending, but use content loss to preserve the original foreground content. Therefore, they strike a balance between preserving the foreground content and smoothing the boundary. However, some smoothed boundary regions are still not satisfactory. MLF [246] can obtain visually appealing results in some cases. Nonetheless, it may erase detailed information (e.g., the small leaves of a pineapple) and fail in handling transparent foreground objects (e.g., plastic bag).
For practical usage, traditional image blending methods are recommended under specific circumstances. For example, when the foreground mask is accurate and there is no high demand for sharp boundary, alpha blending [152] could be used. When the foreground has the same target color as surrounding background, Poisson blending [150] could be used. In other complex cases, learnable image blending methods [246, 243] are recommended, but they may struggle in dealing with transparent foreground and delicate edges. In the cases of transparent foreground (e.g., glass) and delicate edges (e.g., floating hair strands), we recommend improving the performance of image matting methods.